# CAOTE: KV Cache Selection for LLMs via Attention Output Error-Based Token Eviction

### v5

@@ 1 Introduction
A central challenge in long-context inference is the increased computational and memory overhead, particularly during the prefill and generation phases. This is primarily due to the quadratic complexity of self-attention and the growing memory footprint from storing activations over extended sequences. To mitigate quadratic attention computation, Key-Value (KV) caching has become a standard optimization, enabling faster inference by reusing previously computed key-value states (Pope et al., 2023). However, in long-context settings, the memory consumed by the KV cache can surpass that of the model itself (Zhang et al., 2024b), posing a significant bottleneck—especially for memory-constrained environments.
Recent efforts to manage KV cache memory can be broadly categorized into training-based approaches (Zhang et al., 2023; Xiao et al., 2024) and post-training methods. This work focuses on the latter, which operate under fixed memory budgets and dynamically evict tokens from the cache. Early post-training methods such as H2O (Zhang et al., 2024b) retain tokens with the highest cumulative attention scores, while subsequent approaches refine this strategy using variants of attention-based metrics (Oren et al., 2024; Li et al., 2024; Tang et al., 2024; Qin et al., 2025). These methods exploit the sparsity of attention—where a small subset of tokens disproportionately influence the output—to preserve only the most salient cached entries.
@@
Empirically, CAOTE consistently improves performance across a diverse set of benchmarks, including 16 tasks from LongBench, Needle-in-Haystack retrieval, and perplexity evaluations. When combined with recent eviction strategies (Zhang et al., 2024b; Oren et al., 2024; Li et al., 2024), CAOTE yields substantial gains in accuracy and efficiency.
@@ 2 Background
Token eviction is a popular methodology (Xiao et al., 2024; Oren et al., 2024; Zhang et al., 2024b) for decoder-only transformer inferences that prevents KV-Cache from growing linearly as token generation continues by preventing less important tokens from being cached. This has dual benefits; first, it limits memory consumption for the KV-cache and second, it reduces the computational complexity of the attention mechanism. Here, we consider the case of processing the input prompt block-wise in resource-restricted environments. In this case, token eviction can save memory and computation not only in the generation phase, but also in the prefill phase, leading to a shorter time-to-first-token which is especially beneficial when the input prompt is extremely long.
@@ 3 CAOTE: KV Caching through Attention Output-Based Token Eviction
We first introduce CAOTE in Subsection 3.1 and how to compute eviction error in closed-form. This is followed by a discussion of its meta-property, demonstrating its applicability with other score-based attention methods such as H2O (Zhang et al., 2024b) in Subsection 3.2. Finally, we propose an efficient approximation of CAOTE in Subsection 3.3. The general workflow of CAOTE is illustrated in Fig. 1, highlighting that the modifications to existing token eviction methods are minimal.
@@ 3.2 CAOTE with general score-based eviction methods
We consider H2O (Zhang et al., 2024b), where the scores ($H=[h_{1},\dots,h_{b+1}]$) are based on the sum of previous attention scores, leading to $\Sigma_{j=1}^{b+1}h_{j}>1$ during generation-phase as proved in Theorem 36 in Section D.1. In this case, simply dividing each token score by the sum of all scores maps the scores to the range $[0,1]$ and ensures that new scores follow $\sum_{i=1}^{b+1}h_{i}^{\text{norm}}=1$.
@@ 4 Results
In this section, we demonstrate the efficacy of CAOTE for boosting performance on state-of-the-art token eviction methods on a wide range of downstream benchmarks.
@@ 4.1 Experiment Setup
We study the impact of CAOTE on different token eviction methods by evaluating on LongBench (Bai et al., 2024), covering single QA, multiple QA, single/multi-document summarization, synthetic, and code generation tasks. We measure long-context perplexity on the Booksum dataset (Kryściński et al., 2021), and lastly, measure recall accuracy on Needle In A Haystack task (Liu et al., 2024; Kamradt, 2023).
@@
We compare the performance of CAOTE to various token eviction methods including: H2O (Zhang et al., 2024b), TOVA (Oren et al., 2024), and SnapKV (Li et al., 2024), on LLAMA3 models: Llama 3.2-3B-Instruct and Llama 3.1-8B-Instruct (Dubey et al., 2024), and QWEN2.5 models: Qwen 2.5-3B-Instruct and Qwen 2.5-7B-Instruct (Yang et al., 2024) for all subsequent experiments.
@@ 4.2 LongBench
We present the accuracy of Llama 3.1-8B-Instruct, Llama 3.2-3B-Instruct and, Qwen 2.5-3B-Instruct, Qwen 2.5-7B-Instruct using baseline eviction methods with budget of 22k, 44k, both with and without CAOTE in Table 2 and Table 3. We observe that the best average performance is given by SnapKV-FastCAOTE for the Llama3 models, while for Qwen 2.5 models SnapKV-CAOTE performs the best. H2O shows $30\%$ improvement with CAOTE, while TOVA, SnapKV also show overall improvements, making their average accuracy closer to dense accuracy. Moreover, we obserive that for some QA tasks (Qasper, MF-en, Musique, 2WikiMQA), H2O-(Fast)CAOTE performs best.
Additional results for the $6$ k, $8$ k budget are shown in Table 9, Table 10 for Llama3 and Qwen 2.5 respectively, in Section E.1, which follow a trend similar to the 22k, 44k budgets.
@@ 4.3 Perplexity
We use the Booksum dataset (Kryściński et al., 2021) to measure generation perplexity of different eviction methods for various budgets. In Table 4, we show perplexity gap between a model using a given eviction strategy and that of the model without token eviction with cache budgets of 22k, 44k and 66k. We observe that when CAOTE is applied to existing eviction methods, the perplexity either improves or surpasses the perplexity of the baseline model. TOVA-FastCAOTE, SnapKV-CAOTE, and SnapKV-FastCAOTE perform best for 66k, 44k, 22k budgets, respectively, for Llama 3.1-8B-Instruct; for Llama 3.2-3B-Instruct, TOVA-FastCAOTE performs best with 22k and 44k budgets and SnapKV-FastCAOTE beats other methods using 66k and 88k. Perplexity results for Qwen 2.5 models are shown in Table 11 in Section E.2.
@@ 4.4 Needle In A HayStack
Lastly, we run extensive experiments on Needle-In-A-Haystack benchmark (Liu et al., 2024; Kamradt, 2023) and show quantitative results in Table 5 and visualizations for 66k budget Llama 3.1-8B model in Figure 2. We observe in Table 5 that H2O-FastCAOTE performs best for all budgets with Llama 3.2-3B. When using a budget 44k with Llama 3.1-8B, CAOTE boosted H2O outperforms TOVA, SnapKV as well. H2O-CAOTE performs best for Llama 3.1-8B with budget = {22k, 66k} and H2O-FastCAOTE performs best for44k budget for Llama3.1-8b. The gains in precision are especially high for the $4$ k budget for the Llama 3.1-8B model, with an increase of up to $30-60\%$. We can see in Figure 2 that CAOTE improves the state-of-the-art eviction method and is able to predict beyond their budget constraints. Results for Qwen 2.5 models are shown in Table 12 in Section E.3.
@@ 5 Related Work
At the extreme end of efficient KV cache management, token eviction methods have been extensively studied. Leveraging the sparsity of attention in LLMs (Xiao et al., 2024; Sheng et al., 2023; Chen et al., 2021), these methods determine the importance of KV pairs using (learned) rules and retain the pairs with high scores in the cache to approximate the attention output. StreamingLLM (Xiao et al., 2024) observes an attention sink phenomenon, which states that the first few tokens tend to receive the majority of attention weights. To exploit this, it proposes SinkAttention, which prioritizes keeping the initial tokens in the cache while doing sliding window-based attention. Other methods, such as H2O (Zhang et al., 2024b), TOVA (Oren et al., 2024), SnapKV (Li et al., 2024), and RoCO (Ren & Zhu, 2024), retain tokens with high attention scores in the KV cache with various algorithmic modifications. These include preserving the first or last tokens in addition to those with high attention scores or applying smoothing techniques to the attention scores. While these token eviction methods primarily rely on attention scores to assess token importance, CAOTE introduces an orthogonal scoring metric that estimates the impact of values on approximating attention outputs. This metric can complement existing token importance scoring approaches, enhancing other eviction methods.
Furthermore, other lines of work enable layer-wise budget optimization (Feng et al., 2024) by considering the scores of all heads jointly, and selecting top-K nodes, while others (Ge et al., 2023) consider managing memory by keeping/discarding based on token characteristics with baseline token eviction (H2O). Our proposed method, CAOTE is highly flexible, and can be integrated with both the mentioned methods to achieve additional boost on performance.
@@ D.1 H2O scores
D.2 Relation of CAOTE score (eviction error) to logits
@@
Appendix E Additional Results
@@
E.1 LongBench
@@
LongBench result for 66k and 88K for Llama 3.2-3B-Instruct/3.1-8B-Instruct and Qwen 2.5-3B-Instruct/8B-Instruct are shown in Table Table 9, Table 10 respectively. We also include Sink attention results (Xiao et al., 2024) with budget of 66k and 88k. For Llama 3.1-8B-Instruct, TOVA-FastCAOTE performs best for 66k budget, while SnapKV-FastCAOTE for 88k budget. For Llama 3.2-3B-Instruct SnapKV-FastCAOTE performs best for both 66k and 88k budget. On the other hand, for Qwen 2.5-7B-Instruct, SnapKV-CAOTE performs the best for both 66k and 88k, and for Qwen 2.5-3B-Instruct, SnapKV-COATE performs best for 66k budget, and SnapKV-FastCAOTE performs best for 88k budget. Additionally, note that all baseline token eviction methods achieve boost in accuracy when using CAOTE or FastCAOTE.
E.2 Perplexity
@@ E.3 Needle in a Haystack
E.4 Ablations
Difference in Attention Outputs
Figure 3: Normalized error between attention outputs with and without eviction for Llama 3.2-3B-Instruct, measured on 5 samples from Qasper dataset.
In Figure 3, we check the average change in attention score in different layers for various eviction methods to provide insight into the performance gain realized by CAOTE. We measure the difference between the dense attention-output and eviction method based attention output, both with and without CAOTE. We repeat this for each layer and average the result to verify that CAOTE results in smaller $L_{2}$ distance to dense attention-output. Figure 3 shows the normalized mean squared error between the dense attention output and eviction-based attention output for H2O, TOVA and SnapKV, both with and without CAOTE in dashed-lines and solid-lines respectively. We observe that the CAOTE variant always produces a smaller gap from dense attention output (no eviction) compared to baselines, supporting our claim that CAOTE evicts the token which minimizes deviation from true attention output. Furthermore, we show that discrepancy in attention output due to eviction (eviction error) causes discrepancy in logits in D.2; to emphasize that optimizing for eviction error results in less discrepancy in logits and, therefore, less deviation on downstream task accuracy.
Different Prompt Block Sizes
Table 13: LongBench accuracy for different tasks from each sub-category with varying block-prompt-size using Llama 3.2-3B-Instruct model with 44k budget. $\infty$ implies entire prompt is consumed without token eviction and eviction starts during generation. We faced out-of-memory issues with some inputs. For $\infty$ case, HotPotQA gave out-of-memory on a 80GB A100 GPU and ran only on 114 out of 200 samples.
Task H2O H2O-CAOTE TOVA TOVA-CAOTE SnapKV SnapKV-CAOTE
block-size 32 128 $\infty$ 32 128 $\infty$ 32 128 $\infty$ 32 128 $\infty$ 32 128 $\infty$ 32 128 $\infty$
HotPotQA 20.43 24.49 35.71 32.37 40.28 46.37 44.36 45.55 52.93 44.72 45.25 52.94 45.99 46.7 53.67 44.66 45.64 52.67
Qasper 30.8 31.94 32.58 36.74 40.79 38.93 38.65 38.83 40.45 39.39 39.79 39.90 38.71 39.22 39.80 40.18 39.74 39.69
GovReport 5.52 5.44 7.07 30.10 30.82 31.48 29.32 29.41 31.25 29.24 29.39 31.15 29.62 29.79 31.82 30.15 30.07 32.42
Trec 46.50 53.00 56.00 59.00 65.5 72.00 62.00 64.00 71.50 64.00 63.00 71.50 64.00 65.50 71.50 67.50 67.50 71.50
LCC 51.59 51.23 51.51 54.06 55.32 54.51 56.23 55.64 56.34 56.18 55.68 56.39 56.24 55.74 56.56 56.27 55.73 56.59
[…]

### v6

@@ 1 Introduction
A central challenge in long-context inference is the increased computational and memory overhead, particularly during the prefill and generation phases. This is primarily due to the quadratic complexity of self-attention and the growing memory footprint from storing activations over extended sequences. To mitigate quadratic attention computation, Key-Value (KV) caching has become a standard optimization, enabling faster inference by reusing previously computed key-value states (Pope et al., 2023). However, in long-context settings, the memory consumed by the KV cache can surpass that of the model itself (Zhang et al., 2024c), posing a significant bottleneck—especially for memory-constrained environments.
Recent efforts to manage KV cache memory can be broadly categorized into training-based approaches (Zhang et al., 2023; Xiao et al., 2024b) and post-training methods. This work focuses on the latter, which operate under fixed memory budgets and dynamically evict tokens from the cache. Early post-training methods such as H2O (Zhang et al., 2024c) retain tokens with the highest cumulative attention scores, while subsequent approaches refine this strategy using variants of attention-based metrics (Oren et al., 2024; Li et al., 2024; Tang et al., 2024; Qin et al., 2025). These methods exploit the sparsity of attention—where a small subset of tokens disproportionately influence the output—to preserve only the most salient cached entries.
@@
Empirically, CAOTE consistently improves performance across a diverse set of benchmarks, including 16 tasks from LongBench, Needle-in-Haystack retrieval, and perplexity evaluations. When combined with recent eviction strategies (Zhang et al., 2024c; Oren et al., 2024; Li et al., 2024), CAOTE yields substantial gains in accuracy and efficiency.
@@ 2 Background
Token eviction is a popular methodology (Xiao et al., 2024b; Oren et al., 2024; Zhang et al., 2024c) for decoder-only transformer inferences that prevents KV-Cache from growing linearly as token generation continues by preventing less important tokens from being cached. This has dual benefits; first, it limits memory consumption for the KV-cache and second, it reduces the computational complexity of the attention mechanism. Here, we consider the case of processing the input prompt block-wise in resource-restricted environments. In this case, token eviction can save memory and computation not only in the generation phase, but also in the prefill phase, leading to a shorter time-to-first-token which is especially beneficial when the input prompt is extremely long.
@@ 3 CAOTE: KV Caching through Attention Output-Based Token Eviction
We first introduce CAOTE in Subsection 3.1 and how to compute eviction error in closed-form. This is followed by a discussion of its meta-property, demonstrating its applicability with other score-based attention methods such as H2O (Zhang et al., 2024c) in Subsection 3.2. Finally, we propose an efficient approximation of CAOTE in Subsection 3.3. The general workflow of CAOTE is illustrated in Fig. 1, highlighting that the modifications to existing token eviction methods are minimal.
@@ 3.2 CAOTE with general score-based eviction methods
We consider H2O (Zhang et al., 2024c), where the scores ($H=[h_{1},\dots,h_{b+1}]$) are based on the sum of previous attention scores, leading to $\Sigma_{j=1}^{b+1}h_{j}>1$ during generation-phase as proved in Theorem 36 in Section D.1. In this case, simply dividing each token score by the sum of all scores maps the scores to the range $[0,1]$ and ensures that new scores follow $\sum_{i=1}^{b+1}h_{i}^{\text{norm}}=1$.
@@ 4 Results
In this section, we demonstrate the efficacy of CAOTE for boosting performance on state-of-the-art token eviction methods on a wide range of downstream benchmarks on recent frontier models: Llama3 and Qwen2.5.
@@ 4.1 Experiment Setup
We study the impact of CAOTE on different token eviction methods by evaluating on LongBench (Bai et al., 2024), covering single QA, multiple QA, single/multi-document summarization, synthetic, and code generation tasks. We measure long-context perplexity on the Booksum dataset (Kryściński et al., 2021), and lastly, measure recall accuracy on Needle In A Haystack task (Liu et al., 2024; Kamradt, 2023). Details on context and generation lengths are in Appendix Section F.1
@@
We compare the performance of CAOTE to various token eviction methods including: H2O (Zhang et al., 2024c), TOVA (Oren et al., 2024), and SnapKV (Li et al., 2024), on Llama3 models: Llama 3.2-3B-Instruct and Llama 3.1-8B-Instruct (Dubey et al., 2024), and Qwen2.5 models: Qwen 2.5-3B-Instruct and Qwen 2.5-7B-Instruct (Yang et al., 2024) for all subsequent experiments.
@@ 4.2 LongBench
We present the accuracy of Llama 3.1-8B-Instruct, Llama 3.2-3B-Instruct and, Qwen 2.5-3B-Instruct, Qwen 2.5-7B-Instruct using baseline eviction methods with budget of 22k, 44k, both with and without CAOTE in Table 2 and Table 3. We observe that the best average performance is given by SnapKV-FastCAOTE for the Llama3 models, while for Qwen2.5 models SnapKV-CAOTE performs the best. H2O shows $30\%$ improvement with CAOTE, while TOVA, SnapKV also show overall improvements, making their average accuracy closer to dense accuracy. Moreover, we obserive that for some QA tasks (Qasper, MF-en, Musique, 2WikiMQA), H2O-(Fast)CAOTE performs best. We have bolded the best accuracy eviction method for 22k budget in Table 2.
Additional results for the 66k, 88k budget are shown in Table 11, Table 12 for Llama3 and Qwen2.5 respectively, in Section F.2, which follow a trend similar to the 22k, 44k budgets.
@@ 4.3 Perplexity
We use the Booksum dataset (Kryściński et al., 2021) to measure generation perplexity of different eviction methods for various budgets. In Table 4, we show perplexity gap between a model using a given eviction strategy and that of the model without token eviction with cache budgets of 22k, 44k and 66k. We observe that when CAOTE is applied to existing eviction methods, the perplexity either improves or surpasses the perplexity of the baseline model. TOVA-FastCAOTE, SnapKV-CAOTE, and SnapKV-FastCAOTE perform best for 66k, 44k, 22k budgets, respectively, for Llama 3.1-8B-Instruct; for Llama 3.2-3B-Instruct, TOVA-FastCAOTE performs best with 22k and 44k budgets and SnapKV-FastCAOTE beats other methods using 66k and 88k. Perplexity results for Qwen2.5 models are shown in Table 13 in Section F.3.
@@ 4.4 Needle In A HayStack
Lastly, we run extensive experiments on Needle-In-A-Haystack benchmark (Liu et al., 2024; Kamradt, 2023) and show quantitative results in Table 5 and visualizations for 66k budget for Llama3.1-8B-Instruct in Figure 2. We observe in Table 5 that H2O-FastCAOTE performs best for all budgets with Llama3.2-3B-Instruct. When using a budget 44k with Llama3.1-8B-Instruct, CAOTE boosted H2O outperforms TOVA, SnapKV as well. H2O-CAOTE performs best for Llama3.1-8B-Instruct with budget = {22k, 66k} and H2O-FastCAOTE performs best for44k budget for Llama3.1-8b. The gains in precision are especially high for the $4$ k budget for the Llama3.1-8B-Instruct, with an increase of up to $30-60\%$. We can see in Figure 2 that CAOTE improves the state-of-the-art eviction method and is able to predict beyond their budget constraints. Results for Qwen2.5 models are shown in Table 14 in Section F.4.
@@ 5 Related Work
At the extreme end of efficient KV cache management, token eviction methods have been extensively studied. Leveraging the sparsity of attention in LLMs (Xiao et al., 2024b; Sheng et al., 2023; Chen et al., 2021), these methods determine the importance of KV pairs using (learned) rules and retain those with high scores to approximate the attention output.
StreamingLLM (Xiao et al., 2024b) observes an attention sink phenomenon, where the first few tokens receive the majority of attention weights. It proposes SinkAttention, which prioritizes retaining initial tokens while applying sliding window-based attention. As seen in Appendix Table 11, SinkAttention with sliding window-based budgeted key-value cache management performs poor than recent methods with CAOTE. Other methods such as H2O (Zhang et al., 2024c), TOVA (Oren et al., 2024), SnapKV (Li et al., 2024), and RoCO (Ren & Zhu, 2024) retain tokens with high attention scores using various algorithmic strategies — including preserving first/last tokens or applying smoothing to attention scores.
While these approaches primarily rely on attention scores to assess token importance, CAOTE introduces an orthogonal scoring metric that estimates the impact of value vectors on approximating attention outputs. This value-centric perspective complements existing attention-based scoring methods and can be integrated with them to enhance eviction performance.
Recent works such as Quest (Tang et al., 2024) propose using approximate attention scores for efficient cache management. CAOTE can be combined with Quest by approximating value vectors (e.g., via min/max pooling) and computing approximate attention outputs to guide eviction decisions. Similarly, CaM (Zhang et al., 2024b) is based solely on attention scores; CAOTE can enhance this by incorporating value vector information. Notably, both Quest and CaM report results only on LLaMA2, without comparisons to other eviction methods, limiting their relevance to current frontier models.
In contrast, we performed comparison with SnapKV, developed around the same time, demonstrates state-of-the-art performance on modern models like LLaMA3, making it more representative of current deployment scenarios. CAOTE complements SnapKV by introducing a value-centric scoring mechanism that can be integrated with attention-based heuristics to further improve eviction strategies.
Other lines of work focus on layer-wise budget optimization (Feng et al., 2024), selecting top-K nodes across heads, or managing memory based on token characteristics (Ge et al., 2023), often using baseline eviction strategies like H2O. CAOTE is highly flexible and can be integrated with these approaches to achieve further performance gains. Finally, DuoAttention (Xiao et al., 2024a) explore orthogonal directions: addresses attention head selection. This method is complementary to CAOTE, which focuses specifically on token-level eviction.
In summary, while recent methods like SnapKV, Quest, and CaM explore efficient token eviction through attention-based heuristics, CAOTE introduces a fundamentally new direction by leveraging value-centric scoring. Its modular design allows seamless integration with both attention-based and approximate methods, making it highly adaptable across model families — including frontier models like LLaMA3 and Qwen2.5.
@@ D.1 H2O scores
D.2 Relation of CAOTE score to output logits
@@
Appendix E Relation between CAOTE and FastCAOTE
@@
We emperically show in Table 8 that CAOTE and FastCAOTE are correlated, therefore, showing the efficacy of FastCAOTE
Appendix F Additional Results
F.1 Task context and generation lengths
Table 9: Task Input Sizes
Task Type Input Size (tokens)
QA tasks 4k – 16k
Summarization 2.1k – 16k
Few-shot learning 5.2k – 22.4k
Synthetic tasks 7k – 11k
Code 1.5k – 4.2k
Needle-in-the-haystack up to 32k
Table 10: Generation Lengths
Task Type Generation Length (tokens)
QA tasks 100 – 300
Summarization 300 – 800
Multi-document reasoning 500 – 1000
Code $\sim$500
We show the average context and generation lengths of different tasks in Table 9, 10 respectively.
F.2 LongBench
@@
LongBench result for 66k and 88K for Llama 3.2-3B-Instruct/3.1-8B-Instruct and Qwen 2.5-3B-Instruct/8B-Instruct are shown in Table Table 11, Table 12 respectively. We also include Sink attention results (Xiao et al., 2024b) with budget of 66k and 88k. For Llama 3.1-8B-Instruct, TOVA-FastCAOTE performs best for 66k budget, while SnapKV-FastCAOTE for 88k budget. For Llama 3.2-3B-Instruct SnapKV-FastCAOTE performs best for both 66k and 88k budget. On the other hand, for Qwen 2.5-7B-Instruct, SnapKV-CAOTE performs the best for both 66k and 88k, and for Qwen 2.5-3B-Instruct, SnapKV-COATE performs best for 66k budget, and SnapKV-FastCAOTE performs best for 88k budget. Additionally, note that all baseline token eviction methods achieve boost in accuracy when using CAOTE or FastCAOTE.
F.3 Perplexity